### v1

# 1st Dimitrios Bouzoulas Energy and Mechanical Engineering

Abstract
Pedestrian detection in RGB images is a key task in pedestrian safety, as the most common sensor in autonomous vehicles and advanced driver assistance systems is the RGB camera. A challenge in RGB pedestrian detection, that does not appear to have large public datasets, is low-light conditions. As a solution, in this research, we propose an automated infrared-RGB labeling pipeline. The proposed pipeline consists of 1) Infrared detection, where a fine-tuned model for infrared pedestrian detection is used 2) Label transfer process from the infrared detections to their RGB counterparts 3) Training object detection models using the generated labels for low-light RGB pedestrian detection. The research was performed using the KAIST dataset. For the evaluation, object detection models were trained on the generated autolabels and ground truth labels. When compared on a previously unseen image sequence, the results showed that the models trained on generated labels outperformed the ones trained on ground-truth labels in 6 out of 9 cases for the mAP@50 and mAP@50-95 metrics. The source code for this research is available at https://github.com/BouzoulasDimitrios/IR-RGB-Automated-LowLight-Pedestrian-Labeling
Index Terms:
Pedestrian Detection, Low-Light Conditions.
I Introduction
Pedestrian detection has a key role in autonomous vehicles (AVs) and advanced driver assistance systems (ADAS), having the potential to increase safety and lower pedestrian related traffic accidents [1]. According to the World Health Organization (WHO) [2], every year approximately 1.19 million people lose their lives in traffic accidents. The crashes referenced by WHO are a result of a multitude of factors such as poor infrastructure, mechanical problems, and human errors. Looking at the various types of operating conditions nighttime/low-light conditions pose a common and unavoidable threat. This has been verified by the data analysis done in [3] on the data provided by the transportation Statistics and the National Highway Traffic Safety Administration of the United States, where it was found that the probability of a fatal accident during nighttime was nine times higher compared to daytime. These statistics verify the need for systems capable of reliably detecting pedestrians in low-light conditions.
Figure 1: Autolabelling pipeline, using unlabelled low-light IR-RGB image pairs as an input, the system outputs labelled RGB images.
AVs and cars with ADAS currently on the roads and in production have primarily RGB cameras available. Using only RGB cameras can be challenging in edge case conditions such as low-light conditions as RGB model accuracy is highly affected by lighting [4]. A crucial part in developing robust models capable of detecting pedestrians in such conditions is large quantities of labelled data. A large limitation in the labelling process is time.
In this research, we propose an automated Infrared(IR) to RGB labelling pipeline aiming to solve the issues posed by nighttime/low-light conditions, omitting the need for human input. For this goal we leverage the lighting invariance of infrared cameras as the visibility of infrared light is not affected by the lack of visible light. In our approach, we utilise the benefits of infrared images to detect pedestrians in infrared and then transfer the labels to the corresponding RGB image pair. Our proposed approach relies on aligned infrared and RGB images, which can be conveniently collected with a research vehicle equipped with suitable sensor setup. The developed pipeline enables automated labelling with no manual human effort, leaving only data collection as manual work.
The proposed pipeline shows a range of results. As a method the proposed pipeline has not been previously proposed for automated labelling. We allow the reuse and further development by open-sourcing the implementation. The experimental results show that the labels generated from the pipeline can match and in many cases outperform ground truth label performance.
II Related Work
II-A Pedestrian Detection
Aiming to enable AVs and ADAS to detect pedestrians reliably, multiple approaches using various sensors and sensor combinations have been proposed. Multiple camera-based solutions have been proposed over the years [5], targeting use cases with favourable lighting conditions. A recent advancement in purely camera-based pedestrian detection is [6], where a modified YOLOv5 model was proposed resulting in mAP@50 76.9% and mAP@50-95 48.3% mAP scores increase on the WiderPerson dataset [7] and a reduced number of floating point operations making the architecture more suitable for real-time applications. In a similar approach [8] proposed a modified YOLOv8 network achieving an mAP@50 of 54.7% mAP@50-95 of 32.6% and a lower number of floating point operations on the WiderPerson dataset. In [9], an infrared camera was utilised for pedestrian detection. Saliency maps were extracted from infrared images and used as an attention mechanism. The approach was used for enhanced pedestrian detection. The results on the KAIST dataset showed a lowered miss rate of 13.4% for daytime and 19.3% for nighttime compared to the baseline in day and night.
For low-light pedestrian detection, more diverse approaches have been proposed to solve the challenges posed by the poor lighting conditions. To achieve improved focus on pedestrian features [10] proposed a fore-background attention mechanism that enhances pedestrian features while reducing background noise, resulting in improved low-light pedestrian detection. The proposed method scored accurate results across multiple datasets compared to the baseline and other state-of-the-art nighttime pedestrian detection methods.
Sensor fusion approaches are a common solution for pedestrian detection, as they offer increased robustness to a wide variety of conditions, including low-light scenarios. Common sensor fusion-based approaches utilize a RGB-Lidar sensor combination. Following the RGB-Lidar approach [11] proposed a two-branch multilevel pedestrian detection network, and when tested on the KITTI dataset [12] achieved an mAP0.5 of 91.6% and 26.7 fps on the Jetson Orin, making the application both accurate and efficient for real-time applications. Similarly, in [13], using the RGB-Lidar sensor combination and applying a stable diffusion-based approach an improvement of the mAP0.5 metric was achieved of 4.2% for BEVDepth and 4.6% for BEVStereo on the nuScenes dataset.
Similarly to the RGB-Lidar methods a wide range of approaches has been suggested for RGB-Infrared/Thermal sensor combination. With [14] proposing a dual network architecture based on YOLOv5, tested on a wide range of publicly available datasets and achieving mAP@0.5 of 96.6%-INRIA, 89.2% CVC09, 90.5% LSIFIR, 56% FLIR-ADAS, 79.8% CVC14, 72.3% Nightowls, 53.3%-KAIST. Using an SSD-like halfway fusion model network was proposed in [15], where features were extracted from thermal and RGB images with two separate networks before being fused, yielding results for KAIST LLVIP and CVC-14, exceeding state-of-the-art performance for mAP metrcis.
II-B Automatic Labelling
Targeting domain adaptation across different datasets, [16] proposed the usage of pseudo-labels for training pedestrian detection models. The approach used an initial dual-input model trained on a baseline dataset using both RGB and thermal images. This was followed by generating and filtering pseudo labels based on confidence on the target dataset using the pretrained model. The model is then trained on the newly annotated dataset and the entire process is repeated until the model performance becomes stable. The proposed method yielded an All-day mAP@50 score of 80.23% while the same model trained on ground truth labels yielded a mAP@50 86.90% resulting in a performance loss of 6.67% compared to ground truth labels in performance while being fully automated.
To achieve automatic labelling across domains [17] proposed an unsupervised learning transfer framework. The proposed method consists of training a model to perform pedestrian detection, using illumination awareness to generate new data, consequently adapting the network to the new domain, and finally performing the automated labelling task on the new domain. The training and autolabelling happens with three models, a daytime using RGB, a nighttime using thermal and a multimodal one using RGB and thermal. The kept labels are decided by an Illumination estimation network as it estimates which autolabeler performs best based on illumination. This approach achieved the lowest miss rate compared to other methods and state-of-the-art performance on the KAIST and FLIR datasets.
III Methods
This paper presents an automatic pedestrian labelling method for RGB image data, leveraging corresponding IR images (Fig.1). The implementation is aimed at enhancing pedestrian detection in low-light conditions. The proposed architecture is presented in Section III-A. The KAIST dataset used in the paper is described in Section III-B, and the implementation details are presented in Section III-C.
III-A Architecture
First, RGB-IR image pair is taken as input, and pedestrians are detected from the IR image. Then the detections are transferred from the IR frame to the corresponding RGB frame to act as a label. Finally, an RGB prediction model for low-light conditions is trained with the autolabeled data.
III-A1 IR detection
In low-light conditions, pedestrians are usually easier to detect from IR images compared to RGB images. IR images capture the IR radiation emitted from the pedestrians, and the amount of IR radiation emitted is not affected by lighting conditions. As a result, IR images are invariant to lighting conditions, and an IR model trained on daytime data generalizes well to low-light conditions. Additionally, due to the heating effect of the sun, the environment usually emits more IR radiation during the day than at night. Thus, the pedestrian IR footprint stands out from the background better at night. From these premises, we train an IR pedestrian detection model with daytime IR data and use it to detect pedestrians in unseen IR images from low-light conditions. These IR detections are then used as labels for the corresponding RGB frames
III-A2 Label transfer
To train an RGB detection model, the labels must be transferred from the IR frame to the RGB frame. Accurate label transfer requires synchronization and known calibration between the IR and RGB frames. Given a pixel in the IR frame, we need to be able to define its position in the RGB frame. To achieve this, the camera matrices must be known, and the principal points of the IR and RGB cameras must be aligned, or at least one of the cameras must also include depth.
In the KAIST dataset used in this paper, the principal points of the RGB and IR cameras are aligned using a beam splitter, and the frames are rectified to have the same focal length, yielding aligned IR and RGB frames. Thus, we can directly transfer the labels from the IR frame to the RGB frame.
III-A3 Training RGB detection model
The IR detections transferred to the RGB frame are used as autolabels for training a low-light pedestrian detection model for RGB images. An IR camera is needed for autolabelling, but the trained prediction model is fully RGB image-based.
III-B Data
[…]
III-C Evaluation
We chose the YOLOv11[18] for the IR detection task, and trained the m, l, and x variants to compare performance across different model sizes. These are referred to as IR-YOLOv11m, IR-YOLOv11l, and IR-YOLOv11x, respectively. We started with the COCO pretrained weights and fine-tuned the weights for the IR pedestrian detection task using the manually labeled daytime IR images. Training was conducted with the default YOLOv11 hyperparameters for 10 epochs. Labels were then automatically generated for the low-light train samples with each variation. The manual ground truth labels were used as a baseline. Altogether, four different training labels were used in the evaluation:
[…]
We employed YOLOv11 also for the RGB detection task, and again trained the m, l, and x variants. These are referred to as YOLOv11m, YOLOv11l, and YOLOv11,x respectively. We started with the COCO pretrained weights and trained for 10 epochs with the default YOLOv11 hyperparameters. Each variant was trained with each of the training labels, yielding altogether 12 different training label-model combinations.
IV Results
The proposed autolabels were evaluated by comparing the performance of the models trained on autolabels to the models trained on ground truth labels (Table I). Precision, Recall, mAP@50, and mAP@50-95 were chosen as evaluation metrics. The evaluation function presented in [19] was used here. Additionally, examples of detection results are shown in Figure 3.
Table I: Validation test results comparison with green for results above the baseline and red for results below baseline. The precision and recall values have been reported with a confidence threshold of 0.1 and IoU threshold of 0.5.
Labels P R mAP@50 mAP@50-95
Base Model YOLOv11m
IR-YOLOv11m (our) 0.813 0.282 0.306 0.284
IR-YOLOv11l (our) 0.754 0.299 0.311 0.288
IR-YOLOv11x (our) 0.926 0.204 0.313 0.280
KAIST/Ground Truth 0.824 0.269 0.290 0.271
Base Model YOLOv11l
IR-YOLOv11m (our) 0.817 0.306 0.339 0.318
IR-YOLOv11l (our) 0.788 0.325 0.359 0.337
IR-YOLOv11x (our) 0.919 0.204 0.308 0.280
KAIST/Ground Truth 0.822 0.289 0.328 0.306
Base Model YOLOv11x
IR-YOLOv11m (our) 0.894 0.266 0.317 0.294
IR-YOLOv11l (our) 0.896 0.272 0.341 0.317
IR-YOLOv11x (our) 0.935 0.176 0.291 0.265
KAIST/Ground Truth 0.802 0.281 0.318 0.295
In both the mAP@50 and mAP@50-95 metrics, the models trained with the proposed autolabels outperformed the models trained with the ground truth in 6 out of 9 cases. For all prediction model sizes, the highest performance was achieved with some version of our autolabels. Overall, the highest mAP@50 and mAP@50-95 was achieved by YOLOv11l trained on labels generated with IR-YOLOv11l.
The functionality of the proposed method is further verified by the Precision-Recall curves shown in Figure 2. Precision-Recall Curves are presented for YOLOv11m, l and x for generated and ground truth labels with IoU threshold of 0.5. The Precision-Recall curves further verify that the proposed labels match the ground labels in performance and in many cases exceed it as in all cases the Precision-Recall curves follow the same pattern, dictating similar precision recall tradeoffs.
Figure 2: Precission-Recall Curves for YOLOv11m, l and x for Generated and Ground Truth Labels with IoU threshold of 0.5.
Training labels
IR-YOLOv11m IR-YOLOv11l IR-YOLOv11x GT labels GT
Prediction model YOLOv11m Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
YOLOv11l Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
YOLOv11x Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Examples of predictions made with different prediction model - training label combinations. Columns 1-3 correspond to the models trained with our autolabels, column 4 to the models trained with ground truth labels and column 5 is the ground truth.
V Discussion
In this paper, an automated labelling method for low-light pedestrian detection leveraging corresponding IR and RGB images was presented. Different baseline YOLOv11 models were tested for both the IR image-based autolabelling and label evaluation through the training of RGB detection model with the autolabels. Training with the autolabels yielded higher mAP@50 and mAP@50-95 than training with the ground truth labels in 6 out of 9 cases.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: KAIST ground truth edge case labels, where pedestrians are poorly visible in the RGB frame.
The proposed autolabel performance exceeding ground truth labels in some cases can be attributed to edge cases in ground truth labels where pedestrians are extremely poorly visible in the RGB images due to occlusion and poor illumination (Fig. 4). Such samples can cause the model to overfit to the training data, as generalizable detection features are not found.
The proposed automatic labelling requires IR-RGB image pairs, meaning the data collection vehicle must be equipped with an IR camera and an RGB camera, which are calibrated together. Deploying such platforms can be demanding, as IR cameras are expensive and the calibration procedure can be complex. However, the IR-RGB camera setup is only needed for training data generation. Additionally, the proposed automatic labelling method enables the collection of extensive datasets with no manual human labelling. The RGB detection models trained with the autolabelled data only require an RGB camera, which is frequently available in most new cars.
The proposed approach was only verified using the KAIST dataset and further research is needed. The next steps include the gathering of a larger IR-RGB dataset to evaluate the performance consistency in larger datasets as well as the effects of a larger number of edge case scenarios with occlusion and poor illumination. Additionally, state-of-the-art methods in infrared detection should be tested as the proposed pipeline relies solely on infrared detections and their quality.
As the proposed method’s results showed potential for low-light applications, similar approaches can be applied to other low-light tasks. One such task is pose estimation, as it could be implemented in a similar manner to the proposed method. Such research could aid the development of more complex systems such as pedestrian crossing intention predictions enabling more advanced decision-making in low-light conditions. Another promising future application is detecting animals, like deer, in low-light conditions.