### v2

# 1st Dimitrios Bouzoulas    2nd Eerik Alamikkotervo    3rd Risto Ojala

Abstract
Pedestrian detection in RGB images is a key task in pedestrian safety, as the most common sensor in autonomous vehicles and advanced driver assistance systems is the RGB camera. A challenge in RGB pedestrian detection, which does not appear to have many public datasets or efficient autolabelling pipelines, is low-light conditions. As a solution, in this research, we propose an automated infrared-RGB labelling pipeline. The proposed pipeline consists of 1) Infrared detection, where a fine-tuned model for infrared pedestrian detection is used 2) Label transfer process from the infrared detections to their RGB counterparts 3) Training object detection models using the generated labels for low-light RGB pedestrian detection. The research was performed using the KAIST dataset. For evaluation, three object detection models, DETR, YOLO, and RCNN, were trained on the generated autolabels and ground truth labels. When compared on a previously unseen image sequence, the results showed that the models trained on generated labels outperformed the ones trained on ground-truth labels in 5 out of 6 cases for the mAP@50 and LAMR metrics, and outperformed ground-truth on mAP@50-95 in all cases. Acquired results indicate that the proposed autolabelling pipeline could be used for scalable annotation of low-light datasets for pedestrian detection. The source code for this research is available at https://github.com/BouzoulasDimitrios/IR-RGB-autoamed-low-light-pedestrian-labelling
I Introduction
Pedestrian detection has a key role in autonomous vehicles (AVs) and advanced driver assistance systems (ADAS), having the potential to increase safety and lower pedestrian-related traffic accidents [1]. According to the World Health Organization (WHO) [2], every year approximately 1.19 million people lose their lives in traffic accidents. The crashes referenced by WHO are a result of a multitude of factors, such as poor infrastructure, mechanical problems, and human errors. Looking at the various types of operating conditions, nighttime/low-light conditions pose a common and unavoidable threat. This has been verified by the data analysis done in [3] on the data provided by the transportation Statistics and the National Highway Traffic Safety Administration of the United States, where it was found that the probability of a fatal accident during nighttime was nine times higher compared to daytime. These statistics verify the need for systems capable of reliably detecting pedestrians in low-light conditions.
Figure 1: Autolabelling pipeline, using unlabelled low-light IR-RGB image pairs as an input, the system outputs labelled RGB images.
AVs and cars with ADAS currently on the roads and in production have primarily RGB cameras available. Using only RGB cameras can be challenging in low-light conditions, as RGB model accuracy is highly affected by lighting [4]. A crucial part in developing robust models capable of detecting pedestrians in such conditions is large quantities of labelled data. A large limitation in the labelling process is time.
In this research, we propose an automated Infrared (IR) to RGB labelling pipeline aiming to solve the issues posed by nighttime/low-light conditions, omitting the need for human input. For this goal, we leverage the lighting invariance of infrared cameras as visible light has minimal effect on infrared radiation.
In our approach, we utilise the benefits of infrared images to detect pedestrians in infrared and then transfer the labels to the corresponding RGB image. Our proposed approach relies on aligned infrared and RGB images, which can be conveniently collected with a research vehicle equipped with a suitable sensor setup. The developed pipeline enables automated labelling with no manual human effort, leaving only data collection as manual work.
The proposed pipeline has not previously been implemented for autolabelling tasks and shows a range of positive initial results. The autolabels generated using the proposed pipeline performed on par or outperformed the ground truth labels in many cases in key metrics such as map@50, map@50-95, and LAMR.
II Related Work
II-A Pedestrian Detection
Early pedestrian detection approaches utilize hand-crafted features such as Histogram of Oriented Gradients (HOG) [5]. However, currently, the best performance is achieved by deep neural networks that can learn complex detection mechanisms from the data. Originally, convolutional neural network architectures were proposed but recently also transformer-based architectures have gained popularity in the pedestrian detection task, DETR [6] being a popular example. Convolutional architectures are typically further divided into one-stage and two-stage detectors. One-stage detectors predict the bounding box coordinates and related class probabilities in a single forward pass. Common example is the YOLO family [7]. Two-stage detectors first propose regions of interest and then classify and refine the proposals. Common examples are RCNN [8] and its predecessors Fast RCNN [9] and Faster RCNN [10]. To support the development of the pedestrian detection models variety of datasets have been presented over the years, e.g. Caltech Pedestrian Dataset [11], CityPersons dataset [12], KITTI [13] and WiderPerson [14].
II-B Detection in low-light conditions
In low-light conditions, pedestrian detection from conventional visible-spectrum RGB images is significantly more challenging as the lack of visible light leads to low contrast, loss of color information, and severe noise effects. To improve low-light RGB detection, image enhancement methods which aim to restore the low-light image before feeding it to a standard detector, have been proposed [15, 16]. Improved performance can be achieved by optimizing the enhancement for the object detection task specifically by end-to-end training [17, 18]. However, superior performance can be achieved by leveraging additional sensors like a thermal camera or lidar that provide measurements which are much less affected by the lighting conditions. Lidar-camera fusion for object detection has been extensively researched and [19, 20] are notable examples including low-light conditions in the experiments. RGB-thermal fusion has also been explored in multiple studies. A popular baseline solution is provided in [21] and recent work presenting a more advanced fusion strategy in [22]. In [23, 24] teacher-student architecture is proposed where the teacher has access to RGB and thermal images while the students can only rely on RGB images to train a reliable RGB detector for low-light conditions. While most of the available data is from daytime, some datasets addressing the low-light conditions have been published. KAIST [21], LLVIP [25], FLIR-ADAS [26] and CVC-14 [27] provide RGB and thermal images, and Nightowls [28] only provides RGB images.
II-C Automated learning
Detection models perform poorly outside their training domain, and manual annotation is one of the main bottlenecks for training data generation. Thus, research efforts have also focused on ways to reduce or remove the need for manual annotations. A common approach is unsupervised domain adaptation, where the model is adapted from a source domain to a target domain and annotations are only available in the source domain. In [29] RGB-thermal pedestrian detection model was adapted from the KAIST dataset domain to the CVC-14 dataset domain in unsupervised manner with performance close to supervised methods. Also in [30] unsupervised domain adaptation yields comparable results to supervised methods when moving between the CVC-14, KAIST and FLIR ADAS datasets. Another way to reduce the need for labeling is to transfer the labels between sensors. If reliable detections are available in one sensor frame they can be moved to other sensor frames. In [31] bounding boxes from pre-trained RGB detector were transfered to the corresponding thermal frames to train thermal detection models without the need for manual labeling. Arguably for low-light pedestrian detection transfering the labels in opposite direction would be lucrative. Reliable detections on thermal images could be transferred to the corresponding RGB frames to train a RGB prediction model.
III Methods
This paper presents an automatic pedestrian labelling method for RGB image data, leveraging corresponding IR images (Fig.1). The implementation is aimed at enhancing pedestrian detection in low-light conditions. The proposed architecture is presented in Section III-A. The KAIST dataset used in the paper is described in Section III-B, and the implementation details are presented in Section III-C.
III-A Architecture
First, an RGB-IR image pair is taken as input, and pedestrians are detected from the IR image. Then the detections are transferred from the IR frame to the corresponding RGB frame to act as a label. Finally, an RGB prediction model for low-light conditions is trained with the autolabeled data.
III-A1 IR detection
In low-light conditions, pedestrians are usually easier to detect from IR images compared to RGB images. IR images capture the IR radiation emitted from the pedestrians, and the amount of IR radiation emitted by pedestrians is roughly invariant to visible light conditions. As a result, an IR model trained on daytime data generalizes well to low-light conditions. Additionally, due to the heating effect of the sun, the environment usually emits more IR radiation during the day than at night. Thus, the pedestrian IR footprint stands out from the background better at night. From these premises, we train an IR pedestrian detection model with daytime IR data and use it to detect pedestrians in unseen IR images from low-light conditions. These IR detections are then used as labels for the corresponding RGB frames
III-A2 Label transfer
To train an RGB detection model, the labels must be transferred from the IR frame to the RGB frame. Accurate label transfer requires synchronisation and known calibration between the IR and RGB frames. Given a pixel in the IR frame, we need to be able to define its position in the RGB frame. To achieve this, the camera matrices must be known, and the principal points of the IR and RGB cameras must be aligned, or at least one of the cameras must also include depth.
In the KAIST dataset [21] used in this paper, the principal points of the RGB and IR cameras are aligned using a beam splitter, and the frames are rectified to have the same focal length, yielding aligned IR and RGB frames. Thus, we can directly transfer the labels from the IR frame to the RGB frame.
III-A3 Training RGB detection model
The IR detections transferred to the RGB frame are used as autolabels for training a low-light pedestrian detection model for RGB images. An IR camera is needed for autolabelling, but the trained prediction model is fully RGB image-based.
III-B Data
The proposed autolabelling method is evaluated with the KAIST dataset. KAIST includes $\sim$95k samples in total, where each sample includes an aligned IR-RGB frame pair and manual ground truth annotation for the cyclist, person, and people classes. The people class is used if individual persons cannot be distinguished from a group of people. In our experiments, we excluded the samples that include cyclist or people labels, yielding $\sim$91.8k samples. This exclusion was done due to the three classes being too closely related to each other, resulting in extremely poor performance for any models trained on the RGB low-light portion of the dataset. Of the remaining images, $\sim$52.7k samples are from daytime (sequences 00, 01, 02, 06, 07, 08) and $\sim$23.7k samples from lowlight conditions (sequences 03, 04, 05, 09, 10, 11). The IR models were trained with all of the daytime sequences except sequence 02, which was retained for testing, resulting in $\sim$51k training images and $\sim$7.3k testing images. The low-light RGB models were trained with all of the low-light sequences except sequence 09, which was retained for testing, resulting in $\sim$26.7k training images and $\sim$3.5k testing images.
III-C Evaluation
We chose to evaluate our pipeline with the YOLOv11[32], RCNN [10], and DETR[33] architectures. The specific versions chosen were YOLOv11l, Faster RCNN, and RF-DETR. All three models were trained on the daytime IR frames for 20 epochs with the default hyperparameters. These fine-tuned versions of the models are referred to as IR-YOLO, IR-RCNN, and IR-DETR. For all three models, the models were initialized with COCO [34] pretrained weights. The fine-tuned IR models were used for autolabelling the low-light portion of the dataset. To evaluate how the label confidence threshold affects the final results, two label sets were generated with each model, using 0.25 and 0.5 confidence thresholds. Additionally, the models were also trained with ground truth labels, allowing for performance comparisons between models trained on generated and ground truth labels.
Table I: Number of generated autolabels by each IR pedestrian detection model with different confidence thresholds.
Model Confidence Label Count
IR-YOLO 0.25 26712
IR-YOLO 0.5 16388
IR-DETR 0.25 49280
IR-DETR 0.5 29291
IR-RCNN 0.25 21116
IR-RCNN 0.5 19124
Ground Truth - 26779
Same YOLO, DETR, and RCNN variants were also used for the low-light RGB pedestrian detection. Each of the three models was trained on its IR version labels generated with 0.25 and 0.5 confidence thresholds, and the ground truth labels for comparison. This resulted in a total of 9 low-light RGB pedestrian detection models. Similar to the IR models, each model was trained for 20 epochs with the default hyperparameters using COCO pretrained weights as the starting point.
IV Results
The number of labels generated with each of the IR detection models at 0.25 and 0.5 confidence thresholds are presented in Table I. The KAIST ground truth label count is included for comparison. IR-DETR with 0.25 confidence generated the highest number of labels, and IR-YOLO with 0.5 confidence generated the least amount of labels. To assess the accuracy of the generated labels, standard object detection metrics Precision(P), Recall(R), F1, mAP@50,mAP@50-95 and Log Average Miss Rate (LAMR) are also presented at both confidence levels (Table II). IR-DETR had the highest overall performance on the mAP and LAMR metrics, and the labels generated with IR-DETR at 0.5 confidence had the highest F1 score. YOLO had the second-highest and RCNN the lowest performance on the mAP and LAMR metrics. For RCNN and YOLO 0.25 confidence threshold yielded a higher F1 score than a 0.5 confidence threshold. The precision-recall curves for the IR models are presented in Figure 2.
Table II: Autolabelling accuracy metrics for the IR pedestrian detection models with different confidences. Best result bolded in each column.
Conf. 0.5 Conf. 0.25
Model P R F1 P R F1 mAP@50 mAP@50-95 LAMR $\downarrow$
IR-YOLO 0.836 0.370 0.513 0.746 0.544 0.630 0.635 0.245 0.564
IR-RCNN 0.768 0.391 0.519 0.743 0.435 0.549 0.554 0.208 0.623
IR-DETR 0.788 0.528 0.632 0.545 0.741 0.628 0.688 0.284 0.518
Table III: Validation test results comparison with green for results above the baseline and red for results below baseline. The precision and recall values have been reported with an IoU threshold of 0.5.
Labels P R F1 mAP@50 mAP@50-95 LAMR $\downarrow$
Base Model YOLO
IR-YOLO (Conf. 0.5) 0.941 0.171 0.289 0.290 0.128 0.733
IR-YOLO (Conf. 0.25) 0.924 0.236 0.376 0.321 0.138 0.707
KAIST/Ground Truth 0.902 0.196 0.322 0.261 0.116 0.758
Base Model RCNN
IR-RCNN (Conf. 0.5) 0.955 0.202 0.333 0.290 0.125 0.724
IR-RCNN (Conf. 0.25) 0.944 0.221 0.358 0.307 0.127 0.711
KAIST/Ground Truth 0.963 0.207 0.341 0.292 0.115 0.719
Base Model DETR
IR-DETR (Conf. 0.5) 0.809 0.337 0.475 0.437 0.178 0.635
IR-DETR (Conf. 0.25) 0.722 0.369 0.489 0.442 0.183 0.634
KAIST/Ground Truth 0.835 0.315 0.457 0.397 0.154 0.663
The performance of the low-light RGB models trained with the autolabels is evaluated in Table III. The RGB model trained with the manual ground truth annotation acts as the baseline. The models trained with the proposed autolabels outperformed the models trained with the ground truth in 5 out of 6 cases in mAP@50 and LAMR metrics, and in 6 out of 6 cases in the mAP@50-95 metric. For all models, ground truth label performance in mAP and LAMR metrics was exceeded when using 0.25 label confidence, and for YOLO and DETR, also when using 0.5 label confidence.
The best performance overall was achieved by DETR trained on generated labels with a confidence value of 0.25. The precision-recall curves for each RGB model variant are presented in Figure 3. Each architecture exhibits quite similar precision-recall tradeoffs regardless of the training labels used, while the differences between architectures are significantly larger.
Figure 2: Precision-Recall curves for IR models used for autolabelling.
(a) YOLO
(b) RCNN
(c) DETR
Figure 3: Precision-Recall (PR) curves for the YOLO, RCNN, and DETR RGB detection models when trained on autolabels generated with the IR models versus ground truth labels.
Training labels
0.25 Conf 0.50 Conf GT labels GT IR GT RGB
Prediction model
DETR
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
YOLO
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
RCNN
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Examples of predictions made with different prediction model - training label combinations. Columns 1-2 correspond to the models trained with our autolabels, column 3 to the models trained with ground truth labels and columns 4 and 5 are the ground truth in infrared and RGB.
V Discussion
In this paper, an automated labelling method for low-light pedestrian detection leveraging corresponding IR and RGB images was presented. YOLOv11l, RF-DETR, and Faster RCNN architectures were tested for both the IR image-based autolabelling and lowlight RGB detection, trained with the autolabeled data. Training with the autolabels yielded higher mAP@50 and LAMR than training with the ground truth labels in 5 out of 6 cases and higher mAP@50-95 in all cases.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: KAIST ground truth difficult case labels, where pedestrians are poorly visible in the RGB frame.
Training with the autolabels yielded better results than training with the ground truth labels as shown in Table III and demonstrated on the testing data in Figure 4. This difference in performance could be attributed to the hard cases in ground truth labels, where pedestrians are extremely poorly visible in the RGB images due to occlusion and poor illumination, with examples of such cases shown in Figure 5. Such samples can cause the model to overfit to the training data, as generalizable detection features are not found, resulting in lower overall performance. Another possible explanation is the difficulty of the low-light RGB detection task. The mAP scores of the IR models are up to two times higher compared to the RGB models. Likely, the accuracy of the autolabels is not the main bottleneck, but instead the limitations of RGB detection in low-light conditions.
In our experiments, the model architecture was found to play a major role in the final performance. All IR models were trained on the same data, DETR outperforming YOLO and RCNN with a significant margin (Table II). A similar pattern was observed when low-light RGB models were trained on the ground truth, DETR being the best-performing model (Table III). These results underline the importance of model selection, as DETR was able to adapt to both infrared and low-light conditions significantly better using the same data as the other models.
Another parameter explored in this paper was the autolabelling detection confidence threshold. When the confidence threshold was lowered from 0.5 to 0.25, Recall increases and Precision decreases for all models, while the F1 score increases for IR-YOLO and IR-RCNN and maintains a similar value for IR-DETR. (Table II). The low-light RGB models trained on the autolabels generated with a confidence threshold of 0.25 performed the best in all cases (Table III). There seemed to be a correlation between the F1 score of the autolabels and the performance of the resulting RGB prediction model.
Overall, all of the low-light RGB models have a low recall rate, with the highest being 0.369 for DETR trained on the IR-DETR labels with 0.25 confidence. This highlights both the need for more quality training data and the potential upside from sensor fusion approaches. Camera-based detection is highly affected by lighting conditions, while less lighting-sensitive sensors can provide reliable additional sources of information. Finally, this approach does not pose computation time limitations, allowing for the usage of ensemble models for autolabelling.
The proposed approach was only verified using the KAIST dataset, and further research is needed. The next steps include the gathering of a larger IR-RGB dataset to evaluate the performance consistency in larger datasets, as well as the effects of a larger number of edge case scenarios with occlusion and poor illumination. Additionally, state-of-the-art methods in infrared detection should be tested as the proposed pipeline relies solely on infrared detections and their quality.
The proposed automatic labelling requires IR-RGB image pairs, meaning the data collection vehicle must be equipped with an IR camera and an RGB camera, which are calibrated together. Deploying such platforms can be demanding, as IR cameras are expensive and the calibration procedure can be complex. However, the IR-RGB camera setup is only needed for training data generation. Additionally, the proposed automatic labelling method enables the collection of extensive datasets with no manual human labelling. The RGB detection models trained with the autolabelled data only require an RGB camera, which is frequently available in most new cars.
As the proposed method’s results showed potential for low-light applications, similar approaches can be applied to other low-light tasks. One such task is pose estimation, as it could be implemented in a similar manner to the proposed method. Such research could aid the development of more complex systems, such as pedestrian crossing intention prediction, enabling more advanced decision-making in low-light conditions.